rxcklesskaisher:

which of the following values in the set below will make the equation 4x + 2 = 18 true? (Only input the number.) {0, 1, 2, 3, 4}

AZ:

Hey! So we need to solve the equation we have 4x + 2 = 18 To solve it, we want to get 'x' all by itself but to do that, we first need to move 2 to the other side since we're adding 2 to 4x, we need to do the opposite to move it to the other side what is the opposite for addition? is it subtraction? multiplication? division?

Shawnte:

ok so pluck all 5 of those numbers and replace them with x and see which equation is equal to 18

rxcklesskaisher:

multipcation

AZ:

are you sure? What is the opposite of addition? think about how it is in pairs we add and we subtract we multiply or we divide

rxcklesskaisher:

ohhhh i wasnt thinking subtcation

AZ:

there you go! so since we're adding 2 to move it to the other side, we need to subtract 2 on both sides so we have 4x + 2 = 18 we subtract 2 on both sides of the equal sign 4x + 2 - 2 = 18 - 2 so what is 2 - 2 = ? and what is 18 - 2 = ?

rxcklesskaisher:

16 0

rxcklesskaisher:

@AZ

AZ:

Good, so now we have 4x + 2 - 2 = 18 - 2 4x = 16 we're multiplying 4 and x so to get rid of the 4, we have to do the opposite of multiplication to 4 so we can get 'x' all by itself what is the opposite of multiplication

rxcklesskaisher:

divison

AZ:

so divide by 4 on both sides what do you get? 4x / 4 will give us just x but what is 16 / 4 =

rxcklesskaisher:

4

AZ:

so 4x = 16 4x/4 = 16/4 x = 4 and ta-da, you're all done!

rxcklesskaisher:

thanks ll i need is to be walked through it

AZ:

no problem!